# Gas contamination and mitigation in a 100 m3 / 10 bar argon TPC with optical readout: a viability study

### v1

@@ Appendix B Useful analytic limits for capped distributors
The solution to $v_{d}$, for an initial upstream velocity $v_{u}$, can be easily obtained from eqs. 10 and 11:
$v_{d}=\frac{\frac{D^{4}}{d^{4}}\frac{1}{C_{d}^{2}}\rho{v_{u}}-\sqrt{\frac{1}{4}C_{r}^{2}\rho^{2}v_{u}^{2}+4\left(\frac{1}{2}\frac{D^{4}}{d^{4}}\frac{1}{C_{d}^{2}}\rho+\frac{1}{4}C_{r}\rho\right)(P_{u}-P_{{}_{\text{TPC}}})}}{2\left(\frac{1}{2}\frac{D^{4}}{d^{4}}\frac{1}{C_{d}^{2}}\rho+\frac{1}{4}C_{r}\rho\right)}.$ (31)
@@
$\frac{1}{4}C_{r}<<\frac{1}{2}\frac{D^{4}}{d^{4}}\frac{1}{C_{d}^{2}},$ (32)
as long as $d<D$, that is fulfilled in the distributor proposed in text by at least one order of magnitude. This allows neglecting the right-term in both the numerator and denominator of eq. 31. The left-term inside the square root is very small, part through the above relation, part because:
$\frac{1}{2}\rho v_{u}^{2}<<P_{u}-P_{{}_{\text{TPC}}}.$ (33)
This is typical of a capped distributor. It has been verified during simulations that the pressure difference always exceeds the velocity term by at least a factor of $10$, and generally much more. Finally, friction losses from eq. 29 are very small due to the smallness of the ratio $\Delta{L}/{D}$ in the cases discussed (generally, 10% at most), so we may disregard the effect too.
@@
$P_{u}-P_{{}_{\text{TPC}}}\simeq\Delta{P}_{\text{hole}}\simeq\frac{1}{2}\frac{D^{4}}{d^{4}}\frac{1}{C_{d}^{2}}\rho(v_{d}-v_{u})^{2},$ (34)
@@
$(v_{d}-v_{u})\cdot\pi\left(\frac{D}{2}\right)^{2}=\Phi_{\text{hole}}=v_{\text{hole}}\cdot\pi\left(\frac{d}{2}\right)^{2}.$ (35)
Under the assumption of uniform discharge, the flow per hole relates to the total flow as $\Phi_{\text{hole}}=\Phi/N_{\text{holes}}$. The analytic expressions given in eq. 14, 15 can be finally obtained:
$\Delta{P}_{\text{hole}}$ $\simeq$ $\frac{1}{2}\frac{1}{C_{d}^{2}}\rho v_{\text{hole}}^{2},$ (36)
$v_{\text{hole}}$ $\simeq$ Φtotal hole area ($\frac{\Phi}{\textnormal{total hole area (t.h.a.)}}.$ (37)
Although $C_{d}$ is fluid-dependent, in the pressure-dominated regime defined by eq. 33 it approaches the asymptotic limit discussed in text: $C_{d}\simeq 0.63$.
@@
$\theta=\gamma\cdot\arctan\left(\sqrt{\frac{\frac{1}{2}\rho v_{u}^{2}}{\Delta{P}_{\text{hole}}}}\right).$ (38)
@@
$v_{u}=\frac{\Phi}{N_{\text{pipes}}}\frac{1}{\pi(D/2)^{2}},$ (39)
and substituting the expression obtained previously for $\Delta{P}_{\text{hole}}$ in the limit of small angles, we arrive at expression 16 in text:
$\theta_{\text{max}}\simeq\gamma C_{d}\frac{d^{2}}{D^{2}}\frac{L_{\text{dist}}}{\Delta{L}}.$ (40)
@@ Appendix C Time-dependent differential equations
$\frac{dN_{\text{X}}}{dt}=q(t)-\frac{dN_{\text{Ar}}}{dt}\cdot\frac{N_{\text{X}}}{N_{\text{Ar}}},$ (41)
@@ C.1 Constant outgassing rate
To solve the previous differential equation, we first consider the case where the outgassing rate $q(t)$ is constant over time. Under this assumption, we can easily integrate equation (41) obtaining
$N_{\text{X}}(t)=\frac{q}{\frac{1}{N_{\text{Ar}}}\frac{dN_{\text{Ar}}}{dt}}\,\left(1-\exp\left\{-\frac{1}{N_{\text{Ar}}}\,\frac{dN_{\text{Ar}}}{dt}\cdot t\right\}\right),$ (42)
@@
$f_{\text{X}}(t)=\frac{q}{\frac{dN_{\text{Ar}}}{dt}}\,\left(1-\exp\left\{-\frac{1}{N_{\text{Ar}}}\,\frac{dN_{\text{Ar}}}{dt}\cdot t\right\}\right).$ (43)
Expression 42 can be rewritten as a function of the magnitudes used in text if we recall the ideal gas law $P=\frac{N_{\text{Ar}}}{V}\,K\,T$, so $N_{\text{Ar}}=\frac{P\,V}{K\,T}$, that is valid as long as $N_{\text{Ar}}\gg N_{\text{X}}$. Similarly, we obtain:
$\frac{dN_{\text{Ar}}}{dt}=\frac{P}{K\,T}\,\frac{dV}{dt}.$ (44)
And we can re-express eq. 43 as:
$f_{\text{X}}(t)=\frac{Q\cdot A}{P\cdot\Phi}\,\left(1-\exp\left\{-\frac{\Phi}{V}\cdot t\right\}\right),$ (45)
@@ C.2 Power-law outgassing rate
A more realistic model for the outgassing of plastics is a power law, where $q(t)\propto t^{-1/2}$. The general solution of equation (41) is
@@
Assuming again that the concentration is zero at $t=0$ and taking the definition in text, under which $q(t)={Q_{10h}}\cdot\frac{A}{KT}\cdot\sqrt{10/t}$, with $t$ in hours, the previous expression can be written as
$N_{\text{X}}(t)=\exp\left\{-\frac{dN_{\text{Ar}}}{dt}\cdot\frac{1}{N_{\text{Ar}}}\cdot t\right\}\,\frac{\sqrt{10\pi}\cdot Q_{10\text{h}}\cdot\frac{A}{KT}\cdot\text{erfi}\left(\sqrt{\frac{dN_{\text{Ar}}}{dt}\frac{1}{N_{\text{Ar}}}}\sqrt{t}\right)}{\sqrt{\frac{dN_{\text{Ar}}}{dt}\frac{1}{N_{\text{Ar}}}}},$ (47)
@@
$f_{\text{X}}(t)=\exp\left\{-\frac{dN_{\text{Ar}}}{dt}\cdot\frac{1}{N_{\text{Ar}}}\cdot t\right\}\,\frac{\sqrt{10\pi}\cdot Q_{10\text{h}}\cdot\frac{A}{KT}\cdot\text{erfi}\left(\sqrt{\frac{dN_{\text{Ar}}}{dt}\frac{1}{N_{\text{Ar}}}}\sqrt{t}\right)}{\sqrt{\frac{dN_{\text{Ar}}}{dt}\frac{1}{N_{\text{Ar}}}}\,N_{\text{Ar}}}.$ (48)
@@
$f_{\text{X}}(t)=\exp\left\{-\frac{\Phi}{V}\cdot t\right\}\,\frac{\sqrt{10\pi}\cdot Q_{10\text{h}}\cdot A\cdot\text{erfi}\left(\sqrt{\frac{\Phi}{V}}\cdot\sqrt{t}\right)}{\sqrt{\frac{\Phi}{V}}\,P\cdot V}.$ (49)
@@ C.3 Gas displacement during filling
$\frac{dN_{\text{X}}}{dt}=-\frac{dN_{\text{Ar}}}{dt}\cdot\frac{N_{\text{X}}}{N_{\text{Ar}}},$ (50)
@@
$f_{\text{X}}(t)=\exp\left\{-\frac{\Phi}{V}\cdot t\right\}.$ (51)
@@ C.4 Gas accumulation
If the outgassing element is oblivious to purification (as assumed in text for N2), eq. 41 is replaced by:
$\frac{dN_{\text{X}}}{dt}=q(t).$ (52)

### v2

@@ Appendix B Useful analytic limits for capped distributors
The solution to $v_{d}$, for an initial upstream velocity $v_{u}$, can be easily obtained from eqs. 3.1 and 3.2:
$v_{d}=\frac{\frac{D^{4}}{d^{4}}\frac{1}{C_{d}^{2}}\rho{v_{u}}-\sqrt{\frac{1}{4}C_{r}^{2}\rho^{2}v_{u}^{2}+4\left(\frac{1}{2}\frac{D^{4}}{d^{4}}\frac{1}{C_{d}^{2}}\rho+\frac{1}{4}C_{r}\rho\right)(P_{u}-P_{{}_{\text{TPC}}})}}{2\left(\frac{1}{2}\frac{D^{4}}{d^{4}}\frac{1}{C_{d}^{2}}\rho+\frac{1}{4}C_{r}\rho\right)}.$ (B.1)
@@
$\frac{1}{4}C_{r}<<\frac{1}{2}\frac{D^{4}}{d^{4}}\frac{1}{C_{d}^{2}},$ (B.2)
as long as $d<<D$, which is fulfilled in the distributor proposed in text by at least one order of magnitude. This allows neglecting the right-term in both the numerator and denominator of eq. B.1. The left-term inside the square root is very small, part through the above relation, part because:
$\frac{1}{2}\rho v_{u}^{2}<<P_{u}-P_{{}_{\text{TPC}}},$ (B.3)
which is typical of a capped distributor. [^10] The inequalities given by eq. B.2 and eq. B.3, with the $D/d$ ratios employed in text, cause the positive solution in eq. B.1 to yield $v_{d}>v_{u}$, that is nonphysical. It has been verified during simulations that the pressure difference always exceeds the velocity term by at least a factor of $10$, and generally much more. Finally, friction losses from eq. A.3 are very small due to the smallness of the ratio $\Delta{L}/{D}$ in the cases discussed (generally, 10% at most), so we may disregard the effect too.
@@
$P_{u}-P_{{}_{\text{TPC}}}\simeq\Delta{P}_{\text{hole}}\simeq\frac{1}{2}\frac{D^{4}}{d^{4}}\frac{1}{C_{d}^{2}}\rho(v_{d}-v_{u})^{2},$ (B.4)
@@
$(v_{u}-v_{d})\cdot\pi\left(\frac{D}{2}\right)^{2}=\Phi_{\text{hole}}=v_{\text{hole}}\cdot\pi\left(\frac{d}{2}\right)^{2}.$ (B.5)
Under the assumption of uniform discharge, the flow per hole relates to the total flow as $\Phi_{\text{hole}}=\Phi/N_{\text{holes}}$. The analytic expressions given in eq. 3.5, 3.6 can be finally obtained:
$\Delta{P}_{\text{hole}}$ $\simeq$ $\frac{1}{2}\frac{1}{C_{d}^{2}}\rho v_{\text{hole}}^{2},$ (B.6)
$v_{\text{hole}}$ $\simeq$ Φtotal hole area ($\frac{\Phi}{\textnormal{total hole area (t.h.a.)}}.$ (B.7)
Although $C_{d}$ is fluid-dependent, in the pressure-dominated regime defined by eq. B.3 it approaches the asymptotic limit discussed in text: $C_{d}\simeq 0.63$.
@@
$\theta=\gamma\cdot\arctan\left(\sqrt{\frac{\frac{1}{2}\rho v_{u}^{2}}{\Delta{P}_{\text{hole}}}}\right).$ (B.8)
@@
$v_{u}=\frac{\Phi}{N_{\text{pipes}}}\frac{1}{\pi(D/2)^{2}},$ (B.9)
and substituting the expression obtained previously for $\Delta{P}_{\text{hole}}$ in the limit of small angles, we arrive at expression 3.7 in text:
$\theta_{\text{max}}\simeq\gamma C_{d}\frac{d^{2}}{D^{2}}\frac{L_{\text{dist}}}{\Delta{L}}.$ (B.10)
@@ Appendix C Time-dependent differential equations
$\frac{dN_{\text{X}}}{dt}=q(t)-\frac{dN_{\text{Ar}}}{dt}\cdot\frac{N_{\text{X}}}{N_{\text{Ar}}},$ (C.1)
@@ C.1 Constant outgassing rate
To solve the previous differential equation, we first consider the case where the outgassing rate $q(t)$ is constant over time. Under this assumption, we can easily integrate equation (C.1) obtaining
$N_{\text{X}}(t)=\frac{q}{\frac{1}{N_{\text{Ar}}}\frac{dN_{\text{Ar}}}{dt}}\,\left(1-\exp\left\{-\frac{1}{N_{\text{Ar}}}\,\frac{dN_{\text{Ar}}}{dt}\cdot t\right\}\right),$ (C.2)
@@
$f_{\text{X}}(t)=\frac{q}{\frac{dN_{\text{Ar}}}{dt}}\,\left(1-\exp\left\{-\frac{1}{N_{\text{Ar}}}\,\frac{dN_{\text{Ar}}}{dt}\cdot t\right\}\right).$ (C.3)
Expression C.2 can be rewritten as a function of the magnitudes used in text if we recall the ideal gas law $P=\frac{N_{\text{Ar}}}{V}\,K\,T$, so $N_{\text{Ar}}=\frac{P\,V}{K\,T}$, that is valid as long as $N_{\text{Ar}}\gg N_{\text{X}}$. Similarly, we obtain:
$\frac{dN_{\text{Ar}}}{dt}=\frac{P}{K\,T}\,\frac{dV}{dt}.$ (C.4)
And we can re-express eq. C.3 as:
$f_{\text{X}}(t)=\frac{Q\cdot A}{P\cdot\Phi}\,\left(1-\exp\left\{-\frac{\Phi}{V}\cdot t\right\}\right),$ (C.5)
@@ C.2 Power-law outgassing rate
A more realistic model for the outgassing of plastics is a power law, where $q(t)\propto t^{-1/2}$. The general solution of equation (C.1) is
@@
Assuming again that the concentration is zero at $t=0$ and taking the definition in text, under which $q(t)={Q_{\text{10h}}}\cdot\frac{A}{KT}\cdot\sqrt{10/t}$, with $t$ in hours, the previous expression can be written as
$N_{\text{X}}(t)=\exp\left\{-\frac{dN_{\text{Ar}}}{dt}\cdot\frac{1}{N_{\text{Ar}}}\cdot t\right\}\,\frac{\sqrt{10\pi}\cdot Q_{10\text{h}}\cdot\frac{A}{KT}\cdot\text{erfi}\left(\sqrt{\frac{dN_{\text{Ar}}}{dt}\frac{1}{N_{\text{Ar}}}}\sqrt{t}\right)}{\sqrt{\frac{dN_{\text{Ar}}}{dt}\frac{1}{N_{\text{Ar}}}}},$ (C.7)
@@
$f_{\text{X}}(t)=\exp\left\{-\frac{dN_{\text{Ar}}}{dt}\cdot\frac{1}{N_{\text{Ar}}}\cdot t\right\}\,\frac{\sqrt{10\pi}\cdot Q_{10\text{h}}\cdot\frac{A}{KT}\cdot\text{erfi}\left(\sqrt{\frac{dN_{\text{Ar}}}{dt}\frac{1}{N_{\text{Ar}}}}\sqrt{t}\right)}{\sqrt{\frac{dN_{\text{Ar}}}{dt}\frac{1}{N_{\text{Ar}}}}\,N_{\text{Ar}}}.$ (C.8)
@@
$f_{\text{X}}(t)=\exp\left\{-\frac{\Phi}{V}\cdot t\right\}\,\frac{\sqrt{10\pi}\cdot Q_{10\text{h}}\cdot A\cdot\text{erfi}\left(\sqrt{\frac{\Phi}{V}\cdot t}\right)}{\sqrt{\frac{\Phi}{V}}\,P\cdot V}.$ (C.9)
($\sqrt{10}$ having units of $[t^{1/2}]$)
@@ C.3 Gas displacement during filling
$\frac{dN_{\text{X}}}{dt}=-\frac{dN_{\text{Ar}}}{dt}\cdot\frac{N_{\text{X}}}{N_{\text{Ar}}},$ (C.10)
@@
$f_{\text{X}}(t)=\exp\left\{-\frac{\Phi}{V}\cdot t\right\}.$ (C.11)
@@ C.4 Gas accumulation
If the outgassing element is oblivious to purification (as assumed in text for N2), eq. C.1 is replaced by:
$\frac{dN_{\text{X}}}{dt}=q(t).$ (C.12)